# Artificial Intelligence for Microbiology and Microbiome Research

### v1

@@ 3.4.4 Colonization outcome prediction
Microbial communities are typically subject to various environmental perturbations, e.g., antibiotic administration and diet, which can impact the balance of the microbial ecosystem and cause or exacerbate disease [224]. Machine learning models can be trained on some observed communities and make predictions for those unobserved communities upon similar perturbations. For instance, MLP has been used to predict the temporal gut community composition of termite perturbed by six different lignocellulose food sources [225]. In addition to predicting the impact of diet change on microbial composition, machine learning methods have also been used to predict the colonization outcomes of exogenous species for complex microbial communities [226]. Those machine learning methods treat the baseline (i.e., pre-invasion) taxonomic profile as inputs and the steady state abundance of the invasive species as output or mathematically, learn the mapping from the baseline taxonomic profile of a community to the steady state abundance of the invading species. Validation of the approach using synthetic data and two commensal gut bacteria species Enterococcus faecium and Akkermansia muciniphila in hundreds of human stool-derived in vitro microbial communities, showed that machine learning models, including random forest, linear regression/logistic regression, and neural ODE can predict not only the binary colonization outcome but also the final abundance of the invading species [226].
@@
3.4.5 Microbial dynamics prediction
A fundamental question in microbial ecology is whether we can predict the temporal behaviors of complex microbial communities. Traditionally, this problem is addressed using system identification or network reconstruction techniques, which assume specific population dynamics described by a set of ordinary differential equations. For example, the classical GLV model in community ecology, which considers pair-wise interactions, can be represented as a directed, signed, and weighted graph, often referred to as an ecological network. Numerous methods have been developed to infer these dynamics and reconstruct the ecological network using temporal or steady-state data [210]. However, this network-based approach typically assumes that inter-species interactions are exclusively pair-wise, which may not reflect the true nature of complex microbial interactions.
Recently, deep learning techniques have been deployed to predict temporal behaviors of microbiomes. For example, in 2022, Baranwal et al. applied LSTM (a classical variant of RNN) to learn from experimental data on temporal dynamics and functions of microbial communities to predict their future behavior and design new communities with desired functions [228]. Using a significant amount of experimental data, they found that this method outperforms the widely used GLV model in community ecology. In 2023, Thompson et al. proposed the Microbiome Recurrent Neural Network (MiRNN) architecture. Inputs to the MiRNN at time step $t-1$ include the state of species abundances, metabolite concentrations, control inputs, and a latent vector that stores information from previous steps and whose dimension determines the flexibility of the model.The output from each MiRNN block is the predicted system state and the latent vector at the next time step $t$. To avoid the physically unrealistic emergence of previously absent species, a constrained feed-forward neural network outputs zero-valued species abundances if species abundances at the previous time step were zero. The authors demonstrated that MiRNN yielded comparable prediction performance to the LSTM model, but with more than a 50,000 fold reduction in the number of model parameters.
These works are of broad interest to those working on microbiome prediction and design to optimize specific target functions. So far, LSTM and MiRNN have been just applied to synthetic communities with 25 diverse and prevalent human gut species and 4 major health-relevant metabolites (acetate, butyrate, lactate, and succinate). Its potential to large systems, e.g., the human gut microbiome, with thousands of species and metabolites would be interesting to explore. The quality of the training data would be crucial.
In addition to methods specifically designed for predicting microbial dynamics, existing methodologies developed for multiple time series forecasting (MTSF) can also be potentially employed. For example, MTSF-DG is a model capable of learning historical relation graphs and predicting future relation graphs to capture dynamic correlations [229]. Evaluating the performance of these general time series prediction methods in the context of microbial dynamics prediction would be very interesting..
3.4.6 Microbiome data simulation and imputation
Often, we need to generate synthetic microbiome data for testing computational methods or imputing missing data points, and there are two primary approaches to achieve this. First, data can be generated from statistical models, such as SparseDOSSA [230], or various population dynamics models using existing software, e.g., miaSim [231]. miaSim is particularly versatile, offering users the ability to simulate data based on specific assumptions and scenarios using four widely recognized population dynamics models: the stochastic logistic model, MacArthur’s consumer-resource model, Hubbell’s neutral model, and the GLV model, along with several of their derivations. Second, generative deep learning techniques, such as generative adversarial networks (GANs), can be employed to create synthetic data. Recent advancements have introduced several GAN-based methods for generating synthetic microbiome data. For example, MB-GAN [232] learns latent spaces from observed microbial abundances and generates simulated abundances based on these learned distributions. DeepBioGen [233]: This model captures visual patterns of sequencing profiles and generates realistic human gut microbiome profiles. Both MB-GAN and DeepBioGen are designed for data augmentation of single time point microbiome datasets. For longitudinal microbiome data imputation, DeepMicroGen offers a robust solution [234]. This method extracts features that incorporate phylogenetic relationships between taxa using CNN. These features are subsequently processed by a bidirectional RNN-based GAN model, which generates imputed values by learning the temporal dependencies between observations at different time points. These advanced methods enhance our ability to generate high-fidelity synthetic microbiome data, crucial for developing and testing new analytical tools in microbiome research.
@@ 3.4.7 Microbial source tracking
Many MST solvers draw inspiration from the analogy between the MST problem and estimating the mixing proportions of conversation topics in a test document. It has been pointed out that this analogy is problematic [239]. In topic modeling [240], a specialized area within NLP, the objective is to uncover the abstract “topics” present in a set of documents, which can be viewed as static or “dead.” In contrast, MST typically involves dynamic, thriving microbial communities where ecological dynamics significantly influence community assembly and their state, that is, the microbial composition. Given these ecological dynamics, a sink community cannot merely be viewed as a convex mixture of known and unknown sources. Indeed, through numerical simulations, analytical calculations, and real data analysis, compelling evidence has been presented that ecological dynamics impose fundamental challenges in community-based MST [239]. Thus, results from current MST solvers require very cautious interpretation.
@@ 3.6 Precision Nutrition
Machine-learning models have shown remarkable accuracy in predicting metabolite profiles from microbial compositions [256, 257, 258]. Furthermore, the intersection of computational biology with nutrition science has led to notable strides in personalized nutrition and food quality prediction [259, 260, 261]. This emerging field focuses on customizing dietary recommendations to individual biological and physiological profiles, aiming to optimize health outcomes. By employing machine learning algorithms and microbiome data analysis, researchers are able to predict individual responses to various foods and diets, marking a significant advancement in the field of precision nutrition.
@@ 3.6.2 Metabolomic profile prediction
Various machine-learning methods have been developed to solve this problem. For example, MelonnPan uses an elastic net linear regression to model the relative abundance of each metabolite using metagenomic features [275]. It simply models each metabolite individually, missing the opportunity to use shared information across metabolomic features to boost prediction performance. Neural encoder-decoder (NED) leverages the constraints of sparsity and non-negative weights for mapping microbiomes to metabolomes [276]. The use of non-negative weights in NED imposes a stringent constraint on the model, which simplifies the model complexity but may limit the learning capacity. MiMeNet (Microbiome-Metabolome Network) is essentially an MLP that models the community metabolome profile using metagenomic taxonomic or functional features obtained from a microbiome sample [257].
Leveraging the state-of-the-art deep-learning method, neural ordinary differential equations (NODE) [219], Wang et al. developed mNODE (metabolomic profile predictor using neural ordinary differential equations) [258]. Since the input dimension (the number of microbial species) is different from the output data (the number of microbial species), mNODE integrates the NODE as a middle module, sandwiched by two densely connected layers to adjust for data dimension variability. The method shows superior performance in both synthetic and real datasets than existing methods. Additionally, mNODE can naturally incorporate dietary information into its analysis of human gut microbiomes, improving metabolomic profile predictions. Its susceptibility analysis uncovers microbe–metabolite interactions, which can be confirmed with both synthetic and real datasets. Overall, these findings highlight mNODE’s effectiveness in exploring the microbiome–diet–metabolome connection and advancing research in precision nutrition.
@@ 3.6.3 Personalized diet recommendation
Despite significant advancements in metabolic modeling and the integration of machine learning techniques for predicting metabolomic profiles, several open questions remain that could drive future research. One such question is to explore whether integrating multi-omics data (combining metagenomic, transcriptomic, and proteomic data) could further refine these predictions. Additionally, reinforcement learning could potentially be leveraged to generate better personalized dietary recommendations.
@@ 3.7.1 Microorganism detection, identification and quantification
AI techniques, especially supervised machine learning algorithms, are widely used to detect, identify, or quantify microorganisms using various types of data from cultured bacteria [14]. Here we briefly discuss how AI techniques are applied across three different data types. (1) Microscopic Images: Deep learning models, particularly CNNs, have been highly effective in analyzing microscopic images of bacterial colonies [284, 285]. By training on labeled images, these models can classify bacterial species based on their shapes, sizes, arrangements, and staining characteristics (e.g., Gram staining). This approach aids in automating bacterial identification in clinical labs and research, improving the speed and accuracy of microbial diagnostics. (2) Spectroscopy Data: Supervised machine learning algorithms are also employed to analyze spectroscopy data, such as mass spectrometry or Raman spectroscopy, to identify microorganisms [286, 287]. For instance, MALDI-TOF (Matrix-Assisted Laser Desorption/Ionization Time-of-Flight) mass spectrometry generates unique protein “fingerprints” for bacterial species. Machine learning models trained on these spectra can quickly and accurately classify species based on their spectral profiles. Raman spectroscopy, which provides molecular fingerprints of samples, also benefits from machine learning algorithms to classify bacterial species or detect specific metabolic or pathogenic profiles. (3) Volatile Organic Compounds (VOCs): Many bacteria emit VOCs as metabolic byproducts, which can serve as unique biomarkers for microbial identification [288]. Gas chromatography-mass spectrometry (GC-MS) or electronic noses (e-noses) are often used to capture these VOCs. Machine learning models trained on VOC patterns can distinguish bacterial species based on their unique VOC profiles. This approach has potential in medical diagnostics, food safety, and environmental monitoring.
@@ 3.7.2 Antimicrobial susceptibility evaluation
The evaluation of antimicrobial susceptibility has evolved significantly, especially with advancements in genomics and AI. Early approaches focused on using well-known antibiotic resistance genes to predict phenotypic susceptibility, achieving good accuracy for pathogens like Staphylococcus aureus, Escherichia coli, and Klebsiella pneumoniae. However, challenges arose with pathogens such as Pseudomonas aeruginosa, where resistance is driven by gene expression changes, leading to less reliable phenotype predictions. AI has emerged as a promising tool to address these limitations, especially when mutational knowledge is incomplete. Combining machine learning with gene expression data has improved predictive accuracy, as seen in recent studies on P. aeruginosa, achieving over 90% accuracy for resistance to meropenem and tobramycin [289]. Nonetheless, predictions for other antibiotics, such as ceftazidime, remain suboptimal. Combining phenotypic and genotypic data has further enhanced accuracy in rapid diagnostics, as demonstrated by Bhattacharyya et al., who achieved 94-99% accuracy in predicting susceptibility profiles for several bacterial species within hours [290]. The use of whole-genome sequencing (WGS) data in machine learning systems has been extended to predict minimal inhibitory concentrations (MICs) and antibiotic susceptibility, with mixed results. For example, prediction accuracy for ciprofloxacin MICs in E. coli remained relatively low compared to other antibiotics [291]. Similar machine learning approaches have been employed for Mycobacterium tuberculosis [292], viral evolution studies [293], and understanding viral resistance [294], showcasing AI’s broad applicability. We emphasize that while AI techniques show great promise in improving antimicrobial susceptibility testing, challenges remain, particularly in achieving consistent accuracy across different pathogens and antibiotic classes.

### v2

@@ 3.4.3 Microbial dynamics prediction
Predicting the colonization outcomes of exogenous species for complex microbial communities can be considered as a special dynamics prediction problem [336]. Machine learning methods treat the baseline (i.e., pre-invasion) taxonomic profile as inputs and the steady state abundance of the invasive species as output or mathematically, learn the mapping from the baseline taxonomic profile of a community to the steady state abundance of the invading species. Validation of the approach using synthetic data and two commensal gut bacteria species Enterococcus faecium and Akkermansia muciniphila in hundreds of human stool-derived in vitro microbial communities, showed that machine learning models, including random forest, linear regression/logistic regression, and neural ODE can predict not only the binary colonization outcome but also the final abundance of the invading species [336].
@@
3.4.4 Control of microbial communities
Multi-species microbial communities are ubiquitous in natural ecosystems. When engineered for biomanufacturing, synthetic co-cultures outperform monocultures by enhancing productivity and reducing metabolic burden through task compartmentalization across subpopulations. Despite these advantages, precise control of multi-species communities in bioreactors remains a major challenge—the primary reason most industrial bioprocessing still relies on monocultures. Recently, deep reinforcement learning (DRL)—reinforcement learning augmented with deep neural networks—has been successfully applied to regulate interacting microbial populations in bioreactors [169]. Reinforcement learning helps agents learn decision-making through trial and error. DRL improves this by using deep learning to extract decisions from unstructured data without manual state space engineering. DRL algorithms can take in very large inputs (e.g., an image of the raw board state and the history of states) and decide what actions to perform to optimize an objective (e.g., winning the game). A famous DRL algorithm is AlphaGo Zero, learning from playing the ancient Chinese game of Go without using any human knowledge [338]. To control the microbial populations in a bioreactor, feedback from a trained DRL agent effectively maintains species at target abundances. Notably, model-free DRL with bang-bang control outperforms traditional proportional-integral controllers under infrequent sampling conditions. Moreover, a robust control policy can be learned in a single 24-hour experiment using just five parallel bioreactors. Finally, DRL can directly optimize bioprocess yield in co-cultures. Crucially, this model-free approach requires no prior knowledge of population interactions, making it broadly applicable. We anticipate that DRL will play a pivotal role in microbial community control.
3.4.5 Microbiome data simulation
In microbiome research, we often need to generate synthetic data for testing computational methods, augment existing data to enhance model training, or impute missing data points for downstream data analyses. There are two primary approaches to achieve these goals: (1) statistical or population dynamics models; (2) deep generative models.
CAMISIM [339] is a microbial community and metagenome simulator. It can model different microbial abundance profiles, multi-sample time series, and differential abundance studies, includes real and simulated strain-level diversity, and generates second- and third-generation sequencing data from taxonomic profiles or de novo. CAMISIM consists of three steps: (1) Design of the microbial community, which includes selection of the community members and their genomes, and assigning them relative abundances; (2) Metagenome sequencing data simulation; and (3) Postprocessing, where the binning and assembly gold standards are produced. TADA [340] uses available microbiome data and a statistical generative model to create new samples distributed around existing samples, taking into account phylogenetic relationships between microbial species. TADA models two types of variations: true variation due to confounding factors or natural biological variation among samples; sampling variation due to the fact that sequencing takes a random (but not necessarily uniformly random) subsample of the true diversity, creating additional variation around the true proportions. SparseDOSSA [341] is a statistical model designed to represent microbial ecological population structures . It models marginal microbial feature abundances using a zero-inflated lognormal distribution, incorporating additional components for absolute cell counts, the sequence read generation process, and interactions between microbes and their environment. AugCoDa [342] combines key principles from compositional data analysis (e.g., the Aitchison geometry of the simplex and subcompositions) with classical data augmentation techniques (e.g., Mixup [343] and CutMix [344]). PhyloMix [345] leverages the phylogenetic relationships among microbiome taxa as an informative prior to guide the generation of synthetic microbial samples. Leveraging phylogeny, PhyloMix creates new samples by removing a subtree from one sample and combining it with the corresponding subtree from another sample. MIDASim [346] is a fast and simple approach for simulating realistic microbiome data that reproduces the distributional and correlation structure of a template microbiome dataset. MIDASim consists of two steps. The first step generates correlated binary indicators that represent the presence-absence status of all taxa, and the second step generates relative abundances and counts for the taxa that are considered to be present in step 1, utilizing a Gaussian copula to account for the taxon-taxon correlations.
miaSim [347] is a versatile R/Bioconductor package, providing tools to simulate (longitudinal) time series data from popular models in microbial ecology, e.g., Self-organised instability (SOI), Hubbell’s neutral model, generalized Lotka-Volterra (gLV), Ricker model (discrete gLV), Stochastic logistic model, MacArthur’s consumer-resource model (CRM). MIMIC [348] is a Python package designed to advance the simulation, inference, and prediction of microbial community interactions and dynamics. MIMIC integrates a suite of mathematical models, including gLV, Gaussian Processes (GP), and Vector Autoregression (VAR) to offer a versatile framework for analyzing microbial dynamics.
Generative deep learning techniques, such as GAN and AE, can be naturally employed for microbiome data generatation/imputation. For example, MB-GAN [170] learns latent spaces from observed microbial abundances and generates simulated abundances based on these learned distributions. By modifying the discriminator network in the original GAN framework to incorporate microbiome diversity-based measurements, MB-GAN captures visual patterns of sequencing profiles and generates realistic human gut microbiome profiles. DeepBioGen[349] captures local visual patterns of sequencing profiles by training conditional Wasserstein GAN, whose generator and critic networks are composed of up-convolutional and convolutional layers, respectively. DeepBioGen can then generates new sequencing profiles capturing those local visual patterns. phylaGAN [350] is a deep learning framework to augment the existing datasets with generated microbiome data using a combination of conditional GAN and autoencoder. The conditional GAN trains two models against each other to compute larger simulated datasets that are representative of the original dataset, while the autoencoder maps the original and the generated samples onto a common subspace to make the prediction more accurate. mbSparse [171] leverages a feature autoencoder (AE) to learn sample representations and a conditional VAE for data reconstruction. The feature AE generates representative embeddings from the input matrix (e.g., the raw OTU matrix) using a stacked architecture of two dense networks in both the encoder and decoder stages. The conditional VAE enhances the traditional VAE by incorporating conditional variables, enabling the generation of diverse, targeted data tailored to specific conditions. Note that the methods were originally designed for data augmentation of cross-sectional microbiome data. For longitudinal microbiome data imputation, DeepMicroGen [349] offers a robust solution. It extracts features that incorporate phylogenetic relationships between taxa using CNN. These features are subsequently processed by a bidirectional RNN-based GAN model, which generates imputed values by learning the temporal dependencies between observations at different time points.
These approaches enhance our ability to generate high-fidelity synthetic microbiome data, crucial for developing and testing new methods for microbiome data analysis.
@@ 3.4.6 Microbial source tracking
Many MST solvers draw inspiration from the analogy between the MST problem and estimating the mixing proportions of conversation topics in a test document. It has been pointed out that this analogy is problematic [351]. In topic modeling [352], a specialized area within NLP, the objective is to uncover the abstract “topics” present in a set of documents, which can be viewed as static or “dead.” In contrast, MST typically involves dynamic, thriving microbial communities where ecological dynamics significantly influence community assembly and their state, that is, the microbial composition. Given these ecological dynamics, a sink community cannot merely be viewed as a convex mixture of known and unknown sources. Indeed, through numerical simulations, analytical calculations, and real data analysis, compelling evidence has been presented that ecological dynamics impose fundamental challenges in community‐based MST [351]. Thus, results from current MST solvers require very cautious interpretation.
@@ 3.5.2 Retrosynthesis: breaking down a target molecule
3.5.3 Modeling microbial metabolism
Metabolic modeling has long provided critical insights into microbial community metabolism. However, conventional flux balance analysis (FBA) methods typically fail to predict adaptive metabolic and regulatory strategies that promote long-term survival and stability, particularly in heterogeneous communities. Recently, SPAM-DFBA (Self-Playing Microbes in Dynamic FBA) [179] was introduced—an RL framework that reimagines microbial metabolism as a decision-making process. In SPAM-DFBA, each GEM is treated as an agent that learns to regulate metabolic fluxes based on observable environmental states. Through a self-play mechanism, agents evolve adaptive strategies via trial and error, without requiring predefined metabolic or regulatory rules. “Maladaptive” decisions are naturally filtered through simulated selection, enabling the discovery of policies that maximize long-term fitness. Unlike traditional FBA and dynamic FBA (DFBA), which typically assume biomass maximization at a single time point, SPAM-DFBA accounts for the long-term consequences of metabolic decisions in dynamic, multi-species contexts. Greedy biomass optimization often leads to early community extinction—outcomes that RL-trained agents learn to avoid. The resulting flux regulation policies reveal why certain behaviors emerge in real microbial systems and predict stable metabolic strategies in complex, interacting ecosystems with minimal reliance on prior assumptions.
@@ 3.6 Precision Nutrition
Machine-learning models have shown remarkable accuracy in predicting metabolite profiles from microbial compositions [366, 181, 184]. Furthermore, the intersection of computational biology with nutrition science has led to notable strides in personalized nutrition and food quality prediction [367, 368, 185, 369]. This emerging field focuses on customizing dietary recommendations to individual biological and physiological profiles, aiming to optimize health outcomes. By employing machine learning algorithms and microbiome data analysis, researchers are able to predict individual responses to various foods and diets, marking a significant advancement in the field of precision nutrition.
@@ 3.6.2 Metabolomic profile prediction
Various machine-learning methods have been developed to solve this problem. For example, MelonnPan uses an elastic net linear regression to model the relative abundance of each metabolite using metagenomic features [182]. It simply models each metabolite individually, missing the opportunity to use shared information across metabolomic features to boost prediction performance. Similar to how word2vec [382] estimates word probabilities conditioned on a single particular word, mmvec [383] takes microbial sequence counts and metabolite abundances from various samples as the input and outputs the estimated conditional probabilities of observing a metabolite given the presence of a specific microbe. These conditional probabilities generate the entire metabolite abundance profile for a given microbiome sample through a multinomial distribution. Note that in the original application of word2vec, the skip-gram technique (i.e., creating word embeddings that focus on predicting surrounding words based on a specific word or target word) was employed to account for the sequential nature of the text. For microbiome data, there is no clear sequential nature. Therefore, in mmvec, the skip-gram was replaced by multinomial sampling, where a single microbe is randomly sampled from a microbiome sample at each gradient descent step. Neural encoder-decoder (NED) leverages the constraints of sparsity and non-negative weights for mapping microbiomes to metabolomes [183]. The use of non-negative weights in NED imposes a stringent constraint on the model, which simplifies the model complexity but may limit the learning capacity. MiMeNet (Microbiome-Metabolome Network) is essentially an MLP that models the community metabolome profile using metagenomic taxonomic or functional features obtained from a microbiome sample [181].
mNODE [184] leverages a relatively new class of deep neural networks, i.e., neural ODE [329], to predict metabolomic profiles from microbial compositions. Neural ODE can be considered the continuous-depth version of a Residual Network (ResNet). Compared with traditional MLPs, Neural ODEs model continuous-time dynamics, enabling smoother and more interpretable representations of temporal processes. Since the input dimension (the number of microbial species) is typically different from the output data (the number of microbial species), mNODE integrates the neural ODE as a middle module, sandwiched by two densely connected layers to adjust for data dimension variability. mNODE shows superior performance in both synthetic and real datasets than existing methods. Additionally, mNODE can naturally incorporate dietary information into its analysis of human gut microbiomes, further improving metabolomic profile predictions. Its susceptibility analysis can uncover microbe–metabolite interactions, which can be confirmed with both synthetic and real datasets.
@@ 3.6.3 Personalized diet recommendation
Managing blood glucose remains a major challenge for individuals with diabetes, as traditional methods often lack the adaptability required for personalized care. Recently, RL has emerged as a promising approach for optimizing insulin delivery in Type 1 Diabetes Mellitus (T1DM) [391]. The proposed system integrates continuous glucose monitoring (CGM) noise and random carbohydrate intake to emulate real-world variability. A closed-loop simulator was developed using the Identifiable Virtual Patient (IVP) model, to serve as the RL environment. A Proximal Policy Optimization (PPO) agent was trained within this framework, achieving an average Time in Range (TIR) of 73%—demonstrating superior glucose control compared to conventional strategies. This study introduces a personalized, RL-driven insulin therapy solution that adapts dynamically to individual physiology and lifestyle. By leveraging PPO in a realistic simulation environment, it offers a scalable foundation for closed-loop artificial pancreas systems, with significant potential to enhance long-term glycemic outcomes in clinical practice.
Despite significant advancements in metabolic modeling and the integration of machine learning techniques for predicting metabolomic profiles, several open questions remain that could drive future research. One such question is to explore whether integrating multi-omics data (combining metagenomic, transcriptomic, and proteomic data) could further refine these predictions.
@@ 3.7.1 Microorganism detection, identification and quantification
AI techniques, especially supervised machine learning algorithms, are widely used to detect, identify, or quantify microorganisms using various types of data from cultured bacteria [15]. Here we briefly discuss how AI techniques are applied across four different data types. (1) Microscopic Images: Deep learning models, particularly CNNs, have been highly effective in analyzing microscopic images of bacterial colonies [186, 392]. By training on labeled images, these models can classify bacterial species based on their shapes, sizes, arrangements, and staining characteristics (e.g., Gram staining). This approach aids in automating bacterial identification in clinical labs and research, improving the speed and accuracy of microbial diagnostics. (2) Spectroscopy Data: Supervised machine learning algorithms are also employed to analyze spectroscopy data, such as mass spectrometry or Raman spectroscopy, to identify microorganisms [393, 394]. For instance, MALDI-TOF (Matrix-Assisted Laser Desorption/Ionization Time-of-Flight) mass spectrometry generates unique protein “fingerprints” for bacterial species [395]. Machine learning models trained on these spectra can quickly and accurately classify species based on their spectral profiles. Raman spectroscopy, which provides molecular fingerprints of samples, also benefits from machine learning algorithms to classify bacterial species or detect specific metabolic or pathogenic profiles. (3) Volatile Organic Compounds (VOCs): Many bacteria emit VOCs as metabolic byproducts, which can serve as unique biomarkers for microbial identification [396]. Gas chromatography-mass spectrometry (GC-MS) or electronic noses (e-noses) are often used to capture these VOCs. Machine learning models trained on VOC patterns can distinguish bacterial species based on their unique VOC profiles. This approach has potential in medical diagnostics, food safety, and environmental monitoring. (4) Fecal microbial load. The microbiota in individual habitats differ in both relative composition and absolute abundance. While sequencing approaches determine the relative abundances of taxa and genes, they do not provide information on their absolute abundances. Recently, machine learning models, e.g., the eXtreme Gradient Boosting (XGBoost) regression model [397] and the Random Forest model [398] have been applied to predict fecal microbial loads solely from relative abundance data. It has been shown that fecal microbial load is the major determinant of gut microbiome variation and is associated with numerous host factors, including age, diet, and medication [397].
@@ 3.7.2 Antimicrobial susceptibility evaluation
The accurate and rapid determination of antimicrobial susceptibility is crucial for guiding effective therapy and combating the rise of multidrug-resistant pathogens. AI, particularly machine learning, has emerged as a powerful tool to accelerate and improve antimicrobial susceptibility testing (AST) by integrating genotypic, transcriptomic, and phenotypic data.
Early applications of AI focused on predicting phenotypic resistance directly from complex genotypic patterns. For example, MLP with one hidden layer trained on protease mutation profiles achieved high correlation ($R^{2}>0.88$) in predicting lopinavir resistance in human immunodeficiency virus type 1 (HIV-1), revealing previously unrecognized resistance-associated mutations [187].
Machine learning has since been extended to bacterial pathogens. In Escherichia coli, linear regression and feature-selection models trained on whole-genome sequencing (WGS) data accurately predict ciprofloxacin minimum inhibitory concentrations (MICs), with 93% of predictions falling within a four-fold dilution range of the true value when using only four key mutation predictors [399]. Similarly, in Pseudomonas aeruginosa, multimodal classifiers combining gene presence/absence, sequence variation, and transcriptomic profiles achieved very high predictive values (>0.9) for resistance to four commonly administered antimicrobial drugs, demonstrating that gene expression data can further enhance diagnostic accuracy [400].
A major breakthrough has been the fusion of genotypic and phenotypic information in a single rapid assay. GoPhAST-R simultaneously detects key resistance genes and early antibiotic-induced transcriptional changes, then applies a Random Forest classifier to classify susceptibility with 94–99% accuracy [188]. Performed directly from positive blood cultures using a hybridization-based multiplexed RNA detection platform, GoPhAST-R delivers phenotypic-equivalent AST results 24–36 hours faster than conventional culture-based methods, with total assay time under 4 hours.
Machine learning has also proven valuable in refining susceptibility breakpoints. For example, classification and regression tree (CART) analysis of pyrazinamide MIC in Mycobacterium tuberculosis identified a clinical breakpoint of >50 mg/L predictive of treatment failure, aligning with independent hollow-fiber pharmacodynamic simulations and providing an evidence-based threshold for rapid molecular assays [189].
Together, these studies illustrate a clear progression: from purely genotypic prediction of resistance phenotypes, to multimodal integration of genomic and expression data, and ultimately to rapid hybrid genotypic–phenotypic platforms that dramatically shorten time-to-result while maintaining or exceeding the accuracy of gold-standard methods. These AI-enabled approaches promise to transform routine microbiology diagnostics, enabling earlier targeted therapy and improved stewardship of existing antibiotics.